Fermi National Laboratory
 
Operation Summary for March 16, 2012 - March 19, 2012
 

NuMI reported receiving 9.18E18 protons on target for the period from 3/12/12 to 3/19/12.

Friday March 16

The day shift began with Operations monitoring the Linac, Booster, and Main Injector (MI), and with the NuMI, MiniBooNE, SeaQuest, and Meson FTBFT-1017 experiments taking beam.

During the day shift, at 9:28 AM, Linac technicians took 13 minutes and upgraded a Linac RF station’s (LRF1) low-level RF system. Water Group personnel reported at 10:10 AM that they found a water leak on a pond water pump flange at MI-50; Operations passed this on to FESS personnel. A Muon Ring person began a yield study at 10:52 AM using the antiproton target to create secondary particles including pions that decay into muons. Booster personnel reported that they were commissioning a new low level RF phase detector system at 12:25 PM. At 2:19 PM, the NuMI beam permit tripped off when the output for a power supply (E:V118) went low; everything reset. SeaQuest made several accesses during this shift.

During the evening shift, Operations completed a MI tune at 4:13 PM. At 4:28 PM, the NuMI beam permit tripped off for the first of two times in seven and a half hours due to V118; everything reset.

Saturday March 17

The midnight shift began with Operations sending beam to NuMI, MiniBooNE, SeaQuest, and T-1017.

During the midnight shift, Operations reported making MI tune and chromaticity changes at 5:56 AM.

During the day shift, SeaQuest made several accesses during this shift.

During the evening shift, Operations completed a Booster injection tune at 7:24 PM. At 10:49 PM, LRF1 RF and cavity LCW pressures went to zero; Operators filled both. SeaQuest made several accesses during this shift.

Sunday March 18

The midnight shift began with Operations sending beam to NuMI, MiniBooNE, SeaQuest, and T-1017.

During the midnight shift, Operations completed a MI tune and chromaticity change at 12:11 AM. T-1017 accessed this shift.

During the day shift, at 9:35 AM, the H- Source high voltage tripped off; it reset. Operations began a 750 KeV line tune at 11:49 AM. At 1:20 PM, the NuMI beam permit tripped off for the first of two times in about an hour due to V118; everything reset. SeaQuest made two accesses during this shift.

During the evening shift, Operations reported at 4:07 PM that the Booster efficiency dropped due to a change in the line voltage, which also happened around the same time on Saturday.

Monday March 19

The midnight shift began with Operations sending beam to NuMI, MiniBooNE, SeaQuest, and T-1017.

During the midnight shift, Operations adjusted the Booster notch at 2:54 AM and again at 5:57 AM.

Operations Coordinator (OpsCo) Plans for Today

The OpsCo agrees with the plans of Booster, MI, and Muon Ring personnel to conduct studies. SeaQuest continues to commission their beam line.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 03/19/2012    email Fermilab
Security, Privacy, Legal  |  Use of CookiesFermi National Accelerator Laboratory